# Binarization-Aware Adjuster: A Theoretical Framework for Bridging Continuous Optimization and Discrete Inference with Application to Edge Detection

### v2

@@ Abstract
In machine learning, discrete decision-making tasks exhibit a fundamental inconsistency between training and inference: models are optimized using continuous-valued outputs, yet evaluated through discrete predictions. This discrepancy arises from the non-differentiability of discretization operations, weakening the alignment between optimization objectives and practical decision outcomes. To address this, we present a theoretical framework for constructing a Binarization-Aware Adjuster (BAA) that integrates binarization behavior directly into gradient-based learning. Central to the approach is a Distance Weight Function (DWF) that dynamically modulates pixel-wise loss contributions based on prediction correctness and proximity to the decision boundary, thereby emphasizing decision-critical regions while de-emphasizing confidently correct samples. Furthermore, a self-adaptive threshold estimation procedure is introduced to better match optimization dynamics with inference conditions. As one of its applications, we implement experiments on the edge detection (ED) task, which also demonstrate the effectiveness of the proposed method experimentally. Beyond binary decision tasks and ED, the proposed framework provides a general strategy for aligning continuous optimization with discrete evaluation and can be extended to multi-valued decision processes in broader structured prediction problems.
@@
Adjusted Loss Function, Binarization-Aware Adjuster, Distance Weight Function, Continuous-Discrete Alignment, Edge Detection
@@ 1 Introduction
Many real-world vision tasks, such as semantic segmentation, image classification, object recognition, and edge detection (ED) [1, 2, 3, 4, 5], ultimately require discrete decisions (e.g., binary or categorical outputs). However, learning methods, such as deep neural networks, constrained by the need for differentiable computation, are typically optimized using continuous-valued loss functions within gradient-based learning frameworks [6, 7, 8, 9, 10]. As non-differentiable inference operations, such as thresholding, quantization, or category assignment, are excluded from the optimization process, a fundamental gap arises between continuous-valued optimization during training and the discrete inference deployed at test time. This gap may lead to suboptimal performance, as optimization does not always align with decision-critical behavior.
@@
This mismatch is particularly evident in pixel-level binary prediction tasks. For example, edge detection (ED) aims to generate binary edge maps but is typically trained using continuous-valued loss functions such as binary cross-entropy (BCE). Since binarization is non-differentiable, it is excluded from the optimization loop and applied only post hoc during evaluation. As a result, model updates tend to prioritize pixels that are easy to optimize but have limited influence on the binary output, whereas pixels near the decision boundary, which are more critical for final accuracy, receive insufficient emphasis due to their lower weights and higher optimization difficulty.
Several studies have attempted to mitigate the inconsistency between decision behavior and training objectives, typically by changing losses that rebalance decision classes, such as weighted BCE [11] or alternative ones [12, 13]. However, to the best of our knowledge, few existing methods explicitly integrate binarization (or discretization) behavior into gradient-based optimization, leaving the fundamental optimization–inference misalignment unresolved.
@@
To address this issue, we propose a theoretically grounded framework named Binarization-Aware Adjuster (BAA), which explicitly embeds binarization behavior into the optimization process. The core idea is a threshold-aware loss adjustment mechanism based on the proposed Distance Weight Function (DWF), which dynamically reweights pixel-level loss according to prediction correctness and proximity to the decision threshold. This enables the model to allocate greater learning capacity to decision-critical regions while down-weighting confidently correct predictions that exert limited influence on binary outputs. Additionally, a self-adaptive threshold estimation strategy is introduced to align the adjuster with inference behavior. Since the BAA is continuous and differentiable, it can be seamlessly incorporated into existing gradient-based optimization pipelines, allowing discrete inference behavior to influence learning directly.
The proposed framework is designed for binary decision tasks and is applied to the ED task to demonstrate its effectiveness. ED is selected as a representative implementation scenario due to its foundational role in low-level vision and its strong dependence on precise decision boundaries. Furthermore, many structured prediction tasks can be viewed as extensions or generalizations of the ED paradigm. Binary decision tasks are selected since they can display our method most conveniently and clearly. Nevertheless, the proposed BAA mechanism is conceptually extensible to multi-value discrete decision tasks and applicable to a broader range of tasks involving discrete decisions, such as segmentation, classification, and object recognition.
@@
    We identify a fundamental mismatch between continuous optimization and discrete inference in model training and propose the Binarization-Aware Adjuster (BAA) framework to bridge this gap.
@@
    We develop a principled formulation based on the Distance Weight Function (DWF), enabling binarization-aware loss adjustment while retaining full differentiability.
@@
    We apply the proposed framework to the edge detection task and implement extensive experiments on substantial benchmark datasets across multiple architectures.
@@ 2 Previous Works
A large number of learning-based systems perform discrete decision-making while being optimized using continuous-valued loss functions. This discrepancy has motivated studies on improving the alignment between continuous optimization during training and discrete evaluation at inference. In this work, edge detection (ED) is adopted as an illustrative validation scenario due to its strong dependency on binary decisions and its relevance to many structured prediction tasks. In the following, we briefly review related studies with ED as a representative example.
@@ 2.1 Datasets
Historically, early vision research commonly relied on general-purpose datasets. For instance, BSDS300/500 [14] have been used in contour detection, segmentation, and ED. Common datasets such as PASCAL VOC [15], COCO [16], and NYUD [17] have been widely adopted across segmentation, detection, and image classification tasks. With increasing task specialization, dataset selection became more focused, e.g., PASCAL VOC for object detection, NYUD for segmentation, and BSDS500 for contour-based analysis.
For ED specifically, dedicated datasets were developed to provide more accurate edge annotations. BIPED/BIPED2 [18] offers edge annotations for 250 high-resolution campus images, BRIND [6] extended BSDS500 with reannotated edge labels, and UDED [19] provides carefully selected high-quality samples.
Despite these efforts, annotation quality remains a major challenge. Human-labeled edge maps often suffer from inconsistency, ambiguity, and noise. Although robust training strategies and refined labeling protocols have been proposed [20, 21, 22], annotation noise still affects both generalization ability and fair benchmarking.
@@ 2.2 Model Architectures
Deep learning architectures have achieved broad success in computer vision [23], natural language processing [24], and decision systems [25]. Major architectural paradigms, including convolutional neural networks (CNNs), transformer models, and diffusion-based methods, have demonstrated significant cross-domain applicability.
Using ED as an example, the pioneering work HED [11] validated the effectiveness of multi-scale deep supervision using CNNs, outperforming traditional approaches [4, 5] and early feature-based learning methods [26, 27, 28, 29, 30]. Subsequent works, such as RCF [31], BDCN [32], PiDiNet [33], and DexiNet [34], further optimized fusion strategies and model structures. Transformer-based [7, 35] and diffusion-based [8, 9] designs have also been explored. More recently, the Extractor–Selector framework [10] demonstrated improvements through refined feature selection. Among these approaches, CNN-based models remain the most widely adopted due to their robustness and maturity.
@@ 2.3 Loss Functions
Loss function design is critical in model optimization. Binary cross-entropy (BCE) was initially the standard for decision tasks. However, BCE assigns equal importance to positive and negative samples, which is unsuitable for imbalanced data. To address this, Weighted BCE (WBCE) is widely adopted and still the most standard one currently. Other alternatives include Dice loss [36] and tracing loss [37] for geometry-aware optimization, AP-loss [38] and RankED [12] based on ranking formulation, and SWBCE [13] that models uncertainty in manual annotations.
Although these methods improve sample weighting, they generally do not explicitly model the final inference rule, leaving the discretization step separate from the optimization process. Consequently, the fundamental continuous–discrete gap remains unresolved. This limitation motivates the development of our BAA framework.
@@ 2.4 Evaluation Metrics
$F\beta=\frac{(1+\beta^{2})\times\text{Precision}\times\text{Recall}}{\beta^{2}\times\text{Precision}+\text{Recall}},$ (1)
@@
Here, TP denotes true positives, FP false positives, and FN false negatives, with $\beta$ typically set to 1. The most frequently used metrics include:
@@
    ODS (Optimal Dataset Score): Best F1 using a single threshold for the entire dataset.
@@
    OIS (Optimal Image Score): Mean of each image’s best F1 using image-specific thresholds.
For ED, the commonly used metric computation was proposed in [26] with error tolerance to compensate for minor spatial deviations. Recent studies [22] recommend stricter and standardized tolerance settings for more reliable comparison.
@@ 3 Methodology: General Theory
This section presents the theoretical framework of our method. Here, the design for binary decision tasks is proposed for illustrative convenience and simplicity. Note that the discussions can be easily extended to multi-value discrete decision tasks.
@@ 3.1 Motivation and Insight
A key challenge in training binary decision models is the mismatch between training objectives and evaluation protocols. During training, models generate continuous-valued outputs, typically normalized to [0,1], and are optimized using pixel-wise loss functions such as WBCE. In contrast, evaluation metrics are computed on binarized maps, obtained by applying a fixed threshold to these outputs. Since the binarization operation is discontinuous, it cannot be directly incorporated into the training process. This disconnect creates a fundamental issue: the loss function does not explicitly align with the discrete classification decisions that ultimately determine evaluation performance.
This gap introduces a critical inefficiency in optimization. Pixels whose predicted values lie on the correct side and far from the decision threshold, namely those correct pixels with high confidence, tend to dominate the gradient signal since they are easy to optimize. In contrast, pixels with predictions near the threshold, which are more ambiguous and critical to classification, contribute less to the gradients since they are more difficult to optimize. Consequently, the optimizer, driven by gradient magnitude, inadvertently prioritizes ’easy’ but less effective pixels over ’ambiguous’ but crucial ones, due to the nature of the optimization that tends to optimize the less-challenging pixels with higher priority. This misallocation of optimization effort can prevent the optimal parameter from being learned, resulting in suboptimal learning and degraded final performance.
For example, Table 1 reports optimal binarization thresholds for various ED models on multiple datasets under WBCE loss, which typically fall between 0.6 and 0.8. Suppose the binarization threshold is 0.7, then reducing the predicted value of a non-edge pixel from 0.50 to 0.01 has no effect on the final binary edge map, but it significantly reduces the loss. Moreover, such confident-correct predictions are often easy to optimize, which is the reason that they become the correctly predicted ones with high confidence. Therefore, training naturally prioritizes them for their optimization ease since they contribute more to loss, even though further improvement on them yields no benefit for the final decision. In contrast, increasing the predicted value of an edge pixel from 0.69 to 0.71 can be more difficult, which is the reason why its prediction is near the decision threshold. Therefore, though critical for correct classification, it contributes less to the loss and thus is more easily overlooked.
@@
To bridge the continuous optimization and the binarization inference, we propose a novel loss adjustment mechanism, realized by a BAA, that dynamically re-weights pixel-wise loss contributions based on both proximity to the decision threshold and classification correctness. By explicitly incorporating binarization behavior into the training process, the BAA method ensures that the model allocates its learning capacity to the most decision-critical pixels.
@@ 3.2 Adjusted Loss Function Formulation
Although designing the BAA is non-trivial, the idea of using it is conceptually simple. Let us first address the simpler task, formally defining the adjusted loss function in this subsection, and leave the more challenging part, designing the BAA, to the next.
We define an adjusted loss as a loss that incorporates a pixel-wise weighting factor:
$L_{Adj,\delta}(Pred,Gt)=\sum_{i}(Adj(Pred_{i},Gt_{i})+\delta)L(Pred_{i},Gt_{i})$ (3)
where $i$ runs over all pixels, $L$ is a selected baseline pixel-wise loss function such as WBCE [^1] The same symbol $L$ is employed to denote both the loss function on a single element such as a pixel ($L(Pred_{i},Gt_{i})$) and on the whole object such an images ($L(Pred,Gt)$)., PredPred and $Gt$ represent the model prediction and the ground truth, respectively, $Pred_{i}$ and $Gt_{i}$ represent the model prediction and the ground truth of the $i$-th element, respectively, and $\delta\geq 0$ is a pre-defined hyperparameter, directly adding loss $L$ with intensity $\delta$ to the loss to stabilize training. The function $Adj$ is called an Adjuster, where $Adj(Pred_{i},Gt_{i})\in[0,1]$ assigns dynamic weights, emphasizing the training signal according to the pixel’s contribution to the final binary decision.
Despite the simple idea, the problem lies in the design of the adjuster. A naive choice of $Adj(Pred_{i},Gt_{i})$ is the binary mask $\hat{Adj}_{thr}(Pred_{i},Gt_{i})$, defined as:
$\hat{Adj}_{thr}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&0&If\ (Pred_{i}-thr)(Gt_{i}-thr)\geq 0\\ &1\ &Otherwise\end{matrix}\right$. (4)
where thrthr is the binarization threshold. Hence, $L_{Adj,\delta}$ satisfying (providing $\delta=0$):

### v3

# Binarization-Aware Adjuster for Discrete Decision Learning with an Application to Edge Detection

@@ Abstract
Discrete decision tasks in machine learning exhibit a fundamental misalignment between training and inference: models are optimized with continuous-valued outputs but evaluated using discrete predictions. This misalignment arises from the discontinuity of discretization operations, which prevents decision behavior from being directly incorporated into gradient-based optimization. To address this issue, we propose a theoretically grounded framework termed the Binarization-Aware Adjuster (BAA), which embeds binarization characteristics into continuous optimization. The framework is built upon the Distance Weight Function (DWF), which modulates loss contributions according to prediction correctness and proximity to the decision threshold, thereby aligning optimization emphasis with decision-critical regions while remaining compatible with standard learning pipelines. We apply the proposed BAA framework to the edge detection (ED) task, a representative binary decision problem. Experimental results on representative models and datasets show that incorporating BAA into optimization leads to consistent performance improvements, supporting its effectiveness. Overall, this work establishes a principled approach for aligning continuous optimization with discrete decision behavior, with its effectiveness demonstrated in a concrete application setting.
@@
Discrete Decision Learning, Binarization-Aware Adjuster, Distance Weight Function, Continuous-Discrete Alignment, Edge Detection
@@ 1 Introduction
Many real-world vision tasks, such as semantic segmentation, image classification, object recognition, and edge detection (ED) [1, 2, 3, 4, 5], ultimately involve discrete decision processes (e.g., binary or categorical outputs). However, modern learning methods, including deep neural networks, are typically optimized using continuous-valued loss functions within gradient-based learning frameworks due to the requirement of differentiable computation [6, 7, 8, 9, 10]. As non-differentiable inference operations, such as thresholding, quantization, or category assignment, are excluded from the optimization process, a practical inconsistency arises between continuous-valued optimization during training and discrete inference at test time. This inconsistency may lead to suboptimal learning behavior, as the optimization objective does not explicitly reflect decision-critical outcomes.
@@
This issue is particularly evident in pixel-level binary prediction tasks. For example, edge detection (ED) aims to generate binary edge maps but is typically trained using continuous-valued loss functions such as binary cross-entropy (BCE). Since binarization is non-differentiable, it is excluded from the optimization loop and applied only during inference and evaluation. Consequently, model updates tend to emphasize predictions that are easy to optimize but have limited impact on the final binary output, while predictions near the decision boundary, which are more critical for classification accuracy, often receive insufficient attention due to their lower contribution to the loss and higher optimization difficulty.
Several studies have attempted to alleviate the discrepancy between training objectives and decision behavior, typically by modifying loss functions to rebalance decision classes, such as weighted BCE (WBCE) [11], or by adopting alternative loss formulations [12, 13]. While these methods improve model performance, they do not explicitly model the influence of binarization on the optimization process. As a result, the underlying misalignment between continuous optimization and discrete inference remains largely unaddressed.
@@
To address this issue, we propose a theoretically motivated framework termed the Binarization-Aware Adjuster (BAA), which incorporates binarization-related decision characteristics into gradient-based optimization in a principled and continuous manner. The core idea of the BAA is a threshold-aware loss adjustment mechanism based on the proposed Distance Weight Function (DWF), which reweights pixel-level loss contributions according to prediction correctness and proximity to the decision threshold. This formulation is designed to emphasize decision-critical predictions near the binarization boundary while reducing the influence of confidently correct predictions whose further optimization has limited effect on the final discrete outcome. Since the BAA is continuous, it can be directly integrated into standard training pipelines without altering model architectures or inference procedures.
The proposed framework focuses on binary decision learning and is evaluated on the ED task as a representative application. ED is chosen due to its fundamental role in low-level vision and its strong dependence on precise binary decisions at the pixel level, making it well-suited for analyzing the interaction between continuous optimization and discrete inference. While the present study concentrates on ED, the formulation of the BAA is general in nature and provides theoretical insight into how decision boundaries can be more effectively reflected in continuous optimization objectives.
@@
    We analyze a practical inconsistency between continuous-valued optimization and discrete inference in gradient-based learning and formalize its impact on binary prediction tasks.
@@
    We propose the Binarization-Aware Adjuster (BAA), a principled and continuous loss adjustment framework based on the Distance Weight Function (DWF), to incorporate binarization behavior into the optimization process.
@@
    We validate the proposed framework on the edge detection (ED) task through extensive experiments on multiple benchmark datasets and architectures.
@@ 2 Previous Works
A large body of learning-based systems perform discrete decision-making at inference while being optimized using continuous-valued loss functions during training. This discrepancy has motivated extensive research on improving the alignment between continuous optimization objectives and discrete evaluation criteria. In this work, edge detection (ED) is adopted as a representative validation scenario due to its strong dependence on binary decisions and its close connection to a wide range of structured prediction problems. In the following, we review relevant studies using ED as an illustrative example, with a focus on datasets, model architectures, loss functions, and evaluation protocols.
@@ 2.1 Datasets
Early research in computer vision commonly relied on general-purpose datasets. For instance, BSDS300/500 [14] have been widely used in contour detection, segmentation, and ED. Similarly, datasets such as PASCAL VOC [15], COCO [16], and NYUD [17] have supported a variety of tasks including segmentation, detection, and image classification. As research topics became more specialized, dataset construction increasingly focused on task-specific characteristics, e.g., PASCAL VOC for object detection, NYUD for indoor segmentation, and BSDS500 for contour-based analysis.
For ED in particular, several dedicated datasets have been developed to provide more accurate and task-oriented annotations. BIPED and BIPED2 [18] provide high-resolution edge annotations for campus scenes, BRIND [6] extends BSDS500 with refined edge labels, and UDED [19] offers carefully curated samples with improved image and annotation quality.
Despite these efforts, annotation quality remains a persistent challenge. Human-labeled edge maps often exhibit ambiguity, inconsistency, and noise, especially near weak or perceptually subjective boundaries. Although recent studies have proposed robust training strategies and improved annotation protocols [20, 21, 22], annotation noise continues to affect both model generalization and the reliability of quantitative evaluation.
@@ 2.2 Model Architectures
Deep learning architectures have achieved remarkable success across a wide range of domains, including computer vision [23], natural language processing [24], and decision-making systems [25]. Representative architectural paradigms, such as convolutional neural networks (CNNs) and transformers, have demonstrated strong modeling capacity and broad applicability.
In the context of ED, the seminal work HED [11] demonstrated the effectiveness of multi-scale deep supervision within CNN architectures, significantly outperforming traditional edge detectors [4, 5] and earlier feature-based learning approaches [26, 27, 28, 29, 30]. Subsequent CNN-based methods, including RCF [31], BDCN [32], PiDiNet [33], and DexiNet [34], further refined network structures and feature fusion strategies. Transformer-based [7, 35] and diffusion-based [8, 9] approaches have also been explored to capture long-range dependencies and generative priors. More recently, the Extractor–Selector framework [10] introduced a modular design to improve feature selection. Despite architectural diversity, CNN-based models remain the most widely adopted in ED due to their robustness, efficiency, and mature training paradigms.
@@ 2.3 Loss Functions
Loss function design plays a central role in model optimization, particularly for binary decision tasks. BCE was initially the standard choice; however, it assigns equal importance to positive and negative samples, which is suboptimal for the highly imbalanced data distributions commonly encountered, such as in ED. To address this issue, WBCE has become a widely adopted baseline and remains one of the most standard loss formulations in practice.
Beyond WBCE, various alternative loss functions have been proposed to better reflect structural or ranking-based objectives. Examples include Dice loss [36] and tracing loss [37] for geometry-aware optimization, AP-loss [38] and RankED [12] based on ranking formulations, and SWBCE [13], which explicitly models uncertainty in human annotations. While these approaches improve class balance, order cues, or perceptual quality, they generally treat binarization as a post-optimization operation and do not explicitly account for the inference rule during gradient-based learning. Consequently, the continuous–discrete inconsistency persists, motivating the development of the proposed BAA framework.
@@ 2.4 Evaluation Metrics
$F_{\beta}=\frac{(1+\beta^{2})\times\text{Precision}\times\text{Recall}}{\beta^{2}\times\text{Precision}+\text{Recall}},$ (1)
@@
Here, $TP$ denotes true positives, $FP$ false positives, and $FN$ false negatives, with $\beta$ typically set to 1. The most commonly reported metrics include:
@@
    ODS (Optimal Dataset Scale): The best F1 score achieved using a single optimal threshold across the dataset.
@@
    OIS (Optimal Image Scale): The average of image-wise best F1 scores obtained using image-specific optimal thresholds.
For ED, the evaluation protocol with spatial tolerance was originally introduced in [26] to account for localization deviations. More recent studies [22] advocate stricter and standardized tolerance settings to improve the reliability and comparability of quantitative results.
@@ 3 Methodology: General Theory
This section presents the theoretical framework of the proposed method. For clarity, the formulation is developed for binary decision tasks, which serve as a representative and intuitive setting. The discussion naturally extends to multi-valued discrete decision problems without altering the core principles.
@@ 3.1 Motivation and Insight
A fundamental challenge in training binary decision models lies in the inconsistency between training objectives and evaluation protocols. During training, model outputs are continuous-valued, typically normalized to $[0,1]$, and optimized using pixel-wise loss functions such as WBCE. In contrast, evaluation metrics are computed on binarized outputs obtained by applying a fixed threshold. Since binarization is a discontinuous operation, it cannot be directly incorporated into gradient-based optimization. As a result, the loss function does not explicitly reflect the discrete decisions that ultimately determine evaluation performance.
This inconsistency leads to an inefficiency in optimization. Predictions that are already on the correct side of the decision threshold and far from it, namely, confidently correct elements, tend to dominate the gradient signal because they are easy to optimize and contribute substantially to the loss. In contrast, predictions near the decision boundary, which are more ambiguous and critical for correct classification, are harder to optimize and therefore contribute less to the gradients. Consequently, the optimizer is implicitly biased toward improving easy but decision-irrelevant predictions, while allocating insufficient learning capacity to ambiguous yet decision-critical ones. This imbalance can prevent the model from converging to parameters that are optimal with respect to the final binary decisions.
As an illustrative example, Table 1 reports optimal binarization thresholds for several ED models trained with WBCE across multiple datasets. These thresholds typically lie between $0.6$ and $0.8$. Suppose the threshold is set to $0.7$. Reducing the prediction of a non-edge pixel from $0.50$ to $0.01$ can significantly decrease the loss, yet it has no effect on the final binary edge map. Such predictions are often optimized early in training and tend to dominate the optimization process due to their large quantity and easily reducible loss values. In contrast, increasing the prediction of an edge pixel from $0.69$ to $0.71$ directly alters the binary decision outcome, but is generally more difficult to optimize as a pixel with the predicted value close to the decision threshold. Moreover, these pixels typically appear near true edge regions, which are more critical for accurate discrimination, but are relatively sparse in number. As a result, their overall contribution to the loss function is limited, making such decision-critical predictions more likely to be under-emphasized during training.
@@
To address this issue, we propose a loss adjustment mechanism realized by the Binarization-Aware Adjuster (BAA), which dynamically reweights pixel-wise loss contributions based on both classification correctness and proximity to the decision threshold. By explicitly incorporating binarization-related decision characteristics into the optimization process, the proposed approach encourages the model to allocate greater learning capacity to decision-critical elements.
@@ 3.2 Adjusted Loss Function Formulation
We begin by formulating the adjusted loss function, deferring the specific construction of the BAA to the subsequent subsection.
The adjusted loss is defined as:
$L_{Adj,\delta}(Pred,Gt)=\sum_{i}(Adj(Pred_{i},Gt_{i})+\delta)L(Pred_{i},Gt_{i}),$ (3)
where $i$ indexes pixels, $L$ denotes a baseline pixel-wise loss such as WBCE [^1] The same symbol $L$ is used to denote the loss on a single element $L(Pred_{i},Gt_{i})$ and on the entire prediction $L(Pred,Gt)$., PredPred and $Gt$ represent the model prediction and ground truth, respectively, and $\delta\geq 0$ is a constant added to stabilize training. The function $Adj$, referred to as an adjuster, assigns dynamic weights in $[0,1]$ according to each element’s contribution to the final binary decision.
A straightforward but impractical choice of $Adj$ is the hard binary mask:
$\hat{Adj}_{thr}(Pred_{i},Gt_{i})=\left\{\begin{matrix}0&\text{if }(Pred_{i}-thr)(Gt_{i}-thr)\geq 0,\\ 1&\text{otherwise},\end{matrix}\right$. (4)
where thrthr denotes the binarization threshold. This formulation ensures that correctly classified elements incur zero loss, while misclassified ones incur positive loss, directly reflecting the binary decision. However, the resulting loss is discontinuous and unsuitable for gradient-based optimization, motivating the need for a continuous approximation.
3.3 Design of the Binarization-Aware Adjuster
Let thrthr denote the binarization threshold. The adjuster $Adj$ is designed to satisfy the following conditions: